# Classport: Designing Runtime Dependency Introspection for Java

### v1

@@ Abstract.
Runtime introspection of dependencies, i.e., the ability to observe which dependencies are currently used during program execution, is fundamental for Software Supply Chain security. Yet, Java has no support for it. We solve this problem with Classport, a system that embeds dependency information into Java class files, enabling the retrieval of dependency information at runtime. We evaluate Classport on six real-world projects, demonstrating the feasibility in identifying dependencies at runtime. Runtime dependency introspection with Classport opens important avenues for runtime integrity checking.
1. Introduction
Figure 1. Overview of Classport, a novel system that enables runtime dependency introspection in Java.
@@ 1. Introduction
Software systems keep increasing in size and complexity. To speed up implementation and satisfy the growing demand for features, developers often reuse existing libraries (Linskens, 2024). They leverage third-party open-source code, which is integrated into the main application as a dependency. This collection of source code, dependencies, and build tools used to aggregate these two into a final software product is known as the Software Supply Chain (Linskens, 2024).
@@
Using third-party dependencies is essential for productivity and reliability, but it also increases the attack surface (OpenSSF and The Linux Foundation, 2022). Security vulnerabilities that exist in a dependency are inherited by downstream software products, leading to potential compromises. Notable Software Supply Chain incidents include the Log4Shell vulnerability (Turunen, 2021), the Codecov compromise (Sharma, 2021), and the xz catastrophe (Przymus and Durieux, 2025). Assessing and hardening the security of the Software Supply Chain requires increasing its so-called “transparency”, i.e., the visibility on its components at all points in time of the software lifecycle (Okafor et al., 2022).
@@
One of the key tools to support transparency is the production of Software Bill of Materials (SBOM(Cybersecurity and Infrastructure Security Agency, 2024). An SBOM is an exhaustive record of the individual components included in a software system (National Telecommunications and Information Administration, 2021). Maintainers can use this list as input for vulnerability assessment to efficiently detect malicious or vulnerable software packages. Most of the approaches to SBOM operate statically, by looking for dependencies in metadata files, such as manifests or lock files, only at build time (Yu et al., 2024; Rabbi et al., 2024; Cofano et al., 2024). This has two fundamental limitations. First, it does not reflect the actual dependency usage at runtime, i.e., whether a dependency is actually loaded or used in production (Balliu et al., 2023). Even if this does not present a direct risk for the application, it makes maintainers and developers waste a lot of time doing unnecessary work by assessing the security of unreachable dependencies, which slows down the response time to the actual security problems. Second, static analysis of dependencies does not support the maintainers or developers to make runtime decisions based on the actually used dependencies. Basically, the fundamental limitation of SBOMs is that all dependency information is lost at runtime.
@@
Runtime dependency introspection solves these limitations. It is the ability of a piece of software to observe which components are executed during a program’s actual run, including component name and component version. Runtime dependency introspection is the necessary building block for advanced security hardening. First, it is crucial for identifying dependencies that are present in the application but never used at runtime. This simplifies vulnerability management by deprioritizing components that are never used and helps in reducing the attack surface through the removal of unused dependencies. Second, runtime dependency introspection is required to be able to restrict execution privileges at runtime, by mapping the accessed resource to the corresponding dependency (Amusuo et al., 2024).
@@
The Java runtime is known for its good support of object-oriented introspection (aka class, field, and method reflection) (Rigger et al., 2017). Yet, it is completely blind to dependencies: it has no native support for dependency introspection at runtime (Xiao et al., 2025). In Java, the information that uniquely identifies a dependency is present at build time, but it is not preserved at runtime. Heuristics like guessing the origin of dependencies from package names have been implemented in previous works (Amusuo et al., 2024). However, they suffer from fundamental limitations, such as the requirement for dependencies to be compliant with fragile conventions to be correctly identified. Overall, the problem of runtime introspection of dependencies is an open one. To the best of our knowledge, there are currently no principled blueprint for identifying which dependencies are actually executed at runtime in Java.
@@
In this paper, we propose Classport, a novel technique for introspecting dependencies at runtime in Java. In essence, Classport brings dependency information to runtime: supplier, dependency name and version (group id, artifact id and version in Maven lingo). Classport operates in two phases by: (1) embedding dependency information into Java binary artifacts through build-time instrumentation, and (2) retrieving this information at runtime through dynamic instrumentation.
@@
We evaluate Classport on six real-world open-source Java projects. Our experimental protocol measures the ability to embed dependency information into the binary artifacts, and assesses the time and disk space overhead caused by the added information. Then, we test Classport’s ability to identify the correct set of runtime dependencies according to a meaningful workload. We also consider the impact on the running application, incl. the runtime overhead.
@@ 2. Background
Dependency management by package managers and build systems varies across ecosystems, as they handle resolution, versioning, and transitive dependencies differently. In the Java ecosystem, Maven is one of the available build systems (Foundation, 2025). Within Maven, developers declare dependencies in a file called pom.xml. Each of these dependencies is downloaded as a binary artifact called a JAR (The Apache Software Foundation, 2024a), and is uniquely identified by Group ID, Artifact ID, and Version (The Apache Software Foundation, 2024b), this triple being called a GAV coordinate. The Group ID represents the organization or project namespace, e.g., org.apache.commons. The Artifact ID specifies the module or library, e.g., commons-lang3. The Version distinguishes the same library across different releases, e.g., 3.12.0.
At build time, the dependencies of pom.xml are resolved, in order to list the GAV coordinates of both direct and transitive dependencies. This allows Maven to download the exact version to use. In cases of conflicts between multiple versions of the same transitive dependency, Maven’s dependency resolution mechanism selects the version closest to the root of the dependency tree (The Apache Software Foundation, 2024a). The version visible in pom.xml is thus not necessarily the same one that will be used at runtime. In Maven, dependency metadata such as the GAV of dependencies is only available at build time; this information is not preserved at runtime at all (Xiao et al., 2025). The platform makes it impossible to identify the set of dependencies that are actually used at runtime, or to make any runtime decision depending on the dependencies. In other words, Java supports classes, methods, and fields introspection, but not dependency introspection.
@@
3. Design of Classport
@@ 3.1. Embedder
When the Embedder encounters manifest files with a dependency JAR, such as MANIFEST.MF files, it removes signature-related entries, e.g., digest attributes, and copies the remaining content. This is necessary because modifying the contents of a signed JAR, such as adding metadata to class files, invalidates its signature. The Embedder ignores signature files and copies all other files without modification. We refrain from altering non-class files. If these other files are used, they must be referenced by other class files — which are themselves embedded. If they are not used, the Java Virtual Machine will not load them, and hence we don’t need to get provenance later. After processing each dependency JAR, it repackages it into a new JAR. It saves it in a directory, which we refer to as the augmented package repository, as shown in Figure 1.
@@
3.1.1. Embedder as a Maven plugin
@@
Within Classport, we create the custom embed goal. We configure it to fulfill the following needs of the plugin: (1) to have project’s class files available locally. This is fundamental because they have to be embedded. (2) To have dependencies’ class files. Dependencies must be resolved to also have both direct and transitive ones.
The first point is achieved by compiling the target project before executing the plugin. It guarantees that the class files for the project are available for processing. The second point is fulfilled by ensuring that the resolution of dependencies happens before the embedding process [^4] https://maven.apache.org/plugin-tools/apidocs/org/apache/maven/plugins/annotations/ResolutionScope.html.
3.1.2. Information embedded as Java Annotations
@@ 5 String version(); // V
The annotation contains the GAV coordinate, i.e., group, artifact, and version of the dependency to which the class belongs. The Retention annotation (Oracle, 2023b), @Retention(RetentionPolicy.RUNTIME), is used to make the ClassportInfo annotation visible at runtime. After the injection of Software Supply Chain information by the Embedder, the target application is packaged as a uber-JAR.
3.2. Introspector
@@
3.2.1. Introspector as Java agent using the Instrumentation API
@@ 11 Attributes:
1org.apache.pdfbox,pdfbox,3.0.4
2org.apache.pdfbox,fontbox,3.0.4
3org.apache.pdfbox,pdfbox-io,3.0.4
4commons-logging,commons-logging,1.3.4
5info.picocli,picocli,4.7.6
6org.apache.pdfbox,pdfbox-tools,3.0.4
7org.apache.pdfbox,pdfbox-debugger,3.0.4
3.4. Implementation
Classport is implemented in Java. It is made of two modules: the Embedder uses the ASM [^6] https://asm.ow2.io/ library to manipulate the bytecode in the class files, while the Introspector is a Java agent that uses the Instrumentation API (Oracle, 2023a). The implementation is open-source and available on GitHub. [^7] https://github.com/chains-project/classport
4. Experimental Methodology
@@ 4.1. Research Questions
Table 1. Study subjects considered in our experiments.
Project Version Deps. Workload
PDFBox-app 3.0.4 12 Extract text from a PDF file
Certificate-ripper 2.4.1 5 Print the certificate of a website
mcs 0.7.3 4 Lookup dependency coordinates in Maven Central
batik 1.17 6 Convert an SVG to PNG
checkstyle 10.23.0 34 Lint a Java file
zxing 3.5.3 4 Decode 4 QR codes
4.2. Dataset
We aim at evaluating Classport with a dataset of real-world, maintained, open-source Java projects, which use Maven and span a range of application domains, dependency sizes, and usage scenarios. Our dataset includes six projects, with projects varying in complexity, and includes both lightweight tools and large frameworks such as checkstyle. The most active repository is zxing, which has over 30K stars and around 10K forks on GitHub. Table 1 lists each project along with its version, number of dependencies, and the workload we use to exercise the project. The number of dependencies is calculated directly by Maven. The workload is decided by considering a regular use case scenario for each application. Note that we do not assume a minimum coverage. We do not aim to detect all the reachable dependencies, the goal is rather to demonstrate feasibility, that Classport is able to bring dependencies information, i.e., the GAV, at runtime, for a typical use case scenario. We compile all projects for RQ1 and analyze the resulting artifact. For RQ2, we execute all projects to test the Introspector component.
4.3. Methodology of RQ1
@@
Dependency Completeness This property ensures that Classport embeds the uber-JAR with dependency metadata of all dependencies. To assess this, we first run the Embedder on the dataset, generating the annotated uber-JARs. We only consider dependencies that are needed for the application’s execution, excluding test-only dependencies, as they do not end up in the uber-JAR. Then, we analyze the uber-JARs to read the annotations inside the binary artifacts. We consider the embedding process complete in terms of dependency metadata if the dependency metadata, i.e., each Group ID, Artifact ID, and Version, matches the ground truth retrieved by running the Maven command mvn dependency:list on the project under test.
Class Completeness We assess completeness by verifying that the annotation is embedded in each class file of every dependency in the project. Although the dependency can have files other than class files, such as build manifests and various configuration files (Sharma et al., 2025), we only consider class files for annotation as they are the only ones that can be executed. To verify completeness, we first run the Embedder on the study subjects. Then, we analyse the uber-JAR, looking inside each class file. If all Java class files are annotated with the Classport annotation, the embedding process is complete in terms of class metadata.
Build Time Overhead Build time overhead is measured by comparing the time taken to build the uber-JAR with and without the Embedder. In detail, we run package phase of the Maven build process with and without the Embedder and measure the time taken for each run using the Unix time command. We compute the time difference between the two runs and report the overhead as absolute time in seconds.
@@
4.4. Methodology of RQ2
The second research question evaluates the behavior of Classport at runtime, in particular, its capability to introspect runtime dependencies. The objective is to show that 1) Classport can correctly identify which dependencies are currently running and 2) the target application’s runtime behavior is not affected by Classport.
@@ 4.4. Methodology of RQ2
Dynamic Correctness Dynamic correctness is satisfied when the runtime detected dependencies correspond to the ones actually used during the execution of the workload, i.e., the number of false negatives and false positives is equal to zero. First, we define a workload for each application, as shown in Table 1. The workload reflects a regular use case of the selected application. Then, we run the workload on the embedded project, which produces a list of runtime dependencies observed during the execution of the workload. We compare this list with the ground truth, i.e., the list of dependencies resulting from the mvn dependency:list command. If all the dependencies match, this means that Classport can correctly identify dependencies for this workload at runtime. In the case of missing dependencies from the list detected by Classport, this means that either the workload does not trigger the dependencies or that it is a false negative. To assess this, we modify the pom.xml of the project to only include dependencies that are detected by Classport. Specifically, we moved the dependencies not detected by Classport out of the runtime classpath by changing the scopes of dependencies. Finally, we repackage the application and rerun it with the same workload. If the application executes without errors and produces the same output as without Classport, this demonstrates the correctness of Classport in identifying all runtime dependencies that are necessary for the workload.
Overhead To measure the runtime impact of Classport, we use Java Microbenchmarking Harness (JMH) to benchmark the workloads with and without the Classport agent. [^8] https://github.com/openjdk/jmh We prefer microbenchmarking over a simple time difference between executions because microbenchmarking helps recording the time of execution when the JIT compiler has fully optimized a Java application. Microbenchmarking ignores optimization time, which can lead to misleading performance results (Costa et al., 2021). By accounting for Java Virtual Machine (JVM) optimizations such as JIT compilation and warm-up, JMH provides a more accurate assessment of the runtime overhead introduced by Classport.

### v2

@@ Abstract
Runtime introspection of dependencies, i.e., the ability to observe which dependencies are currently used during program execution, is fundamental for Software Supply Chain security. Yet, Java has no support for it. We solve this problem with Classport, a blueprint and system that embeds dependency information into Java class files, enabling the retrieval of dependency information at runtime. We evaluate Classport on six real-world projects, demonstrating the feasibility in identifying dependencies at runtime.
keywords:
Software Supply Chain, Java, Maven
1 Introduction
@@
Software systems keep increasing in size and complexity. To speed up implementation and satisfy the growing demand for features, developers often reuse existing libraries [1]. They leverage third-party open-source code, which is integrated into the main application as a dependency. This collection of source code, dependencies, and build tools used to build a final software product is known as the Software Supply Chain [1].
@@
Using third-party dependencies is essential for productivity and reliability, but it also increases the attack surface [2]. Security vulnerabilities that exist in a dependency are inherited by downstream software products, leading to potential compromises. Notable Software Supply Chain incidents include the Log4Shell vulnerability [3], the Codecov compromise [4], and the xz catastrophe [5]. Assessing and hardening the security of the Software Supply Chain requires increasing its so-called “transparency”, i.e., the visibility on its components at all points in time of the software lifecycle [6].
@@
One of the key tools to support transparency is the production of Software Bill of Materials (SBOM[7]. An SBOM is an exhaustive record of the individual components included in a software system [8]. Maintainers can use this list as input for vulnerability assessment to detect malicious or vulnerable software packages. Most of the approaches to SBOM operate statically, by looking for dependencies in metadata files, such as manifests or lock files, only at build time [9, 10, 11]. This has two fundamental limitations. First, it does not reflect the actual dependency usages at runtime, i.e., whether a dependency is actually loaded or used in production [12]. Even if this does not present a direct risk for the application, it makes maintainers and developers waste a lot of time doing unnecessary work by assessing the security of unreachable dependencies, which slows down the response time to the actual security problems. Second, static analysis of dependencies does not support the maintainers or developers to make runtime decisions based on the actually used dependencies. Basically, the fundamental limitation of SBOMs is that all dependency information is lost at runtime.
@@
Runtime dependency introspection solves these limitations. It is the ability of a piece of software to observe which components are executed during a program’s actual run, including component name and component version. Runtime dependency introspection is the necessary building block for advanced security hardening. First, it is crucial for identifying dependencies that are present in the application but never used at runtime. This simplifies vulnerability management by deprioritizing components that are never used and helps in reducing the attack surface through the removal of unused dependencies. Second, runtime dependency introspection is required to be able to restrict execution privileges at runtime, by mapping the accessed resource to the corresponding dependency [13].
@@
The Java runtime is known for its good support of object-oriented introspection (aka class, field, and method reflection) [14]. Yet, it is completely blind to dependencies: it has no native support for dependency introspection at runtime [15]. In Java, the information that uniquely identifies a dependency is present at build time, but it is not preserved at runtime. Heuristics like guessing the origin of dependencies from package names have been implemented in previous works [13]. However, they suffer from fundamental limitations, such as the requirement for dependencies to be compliant with fragile conventions to be correctly identified. Overall, the problem of runtime introspection of dependencies is an open one. To the best of our knowledge, there are currently no principled architecture for identifying which dependencies are actually executed at runtime in Java.
@@
In this paper, we propose Classport, a novel technique for introspecting dependencies at runtime in Java. In essence, Classport brings dependency information to runtime: supplier, dependency name and version (group id, artifact id and version in Maven lingo). Classport operates in two phases by: 1. embedding dependency information into Java binary artifacts through build-time instrumentation, and 2. retrieving this information at runtime through dynamic instrumentation.
@@
We evaluate Classport on six real-world open-source Java projects. Our experimental protocol measures the ability to embed dependency information into the binary artifacts, and assesses the time and disk space overhead caused by the added information. Then, we test Classport’s ability to correctly identify the set of runtime dependencies using the test suite of the project. We also measure the impact on the running application, incl. the runtime overhead.
@@ 2 Background
Dependency management by package managers and build systems varies across ecosystems, as they handle resolution, versioning, and transitive dependencies differently. In the Java ecosystem, Maven is one of the available build systems [16]. Within Maven, developers declare dependencies in a file called pom.xml. Each of these dependencies is downloaded as a binary artifact called a JAR [17], and is uniquely identified by Group ID, Artifact ID, and Version [18], this triple being called a GAV coordinate. The Group ID represents the organization or project namespace, e.g., org.apache.commons. The Artifact ID specifies the module or library, e.g., commons-lang3. The Version distinguishes the same library across different releases, e.g., 3.12.0.
At build time, the dependencies of pom.xml are resolved, in order to list the GAV coordinates of both direct and transitive dependencies. This allows Maven to download the exact version to use. In cases of conflicts between multiple versions of the same transitive dependency, Maven’s dependency resolution mechanism selects the version closest to the root of the dependency tree [17]. The version visible in pom.xml is thus not necessarily the same one that will be used at runtime. In Maven, dependency metadata such as the GAV of dependencies is only available at build time; this information is not preserved at runtime at all [15]. The platform makes it impossible to identify the set of dependencies that are actually used at runtime, or to make any runtime decision depending on the dependencies. In other words, Java supports classes, methods, and fields introspection, but not dependency introspection.
@@
3 Design of Classport
Figure 1: Overview of Classport, a novel system that enables runtime dependency introspection in Java.
@@ 3.1 Embedder
When the Embedder encounters manifest files with a dependency JAR, such as MANIFEST.MF files, it removes signature-related entries, e.g., digest attributes, and copies the remaining content. This is necessary because modifying the contents of a signed JAR, such as adding metadata to class files, invalidates its signature. However, Classport is designed to be used by application developers and they can resign the JAR with their keys after the embedding process. For instance, consider a project $P$ authored by an application developer $A_{1}$ that depends on a JAR $D$ signed by a third-party library developer $A_{2}$. During embedding, we strip $A_{2}$’s signature from $D$ so that we can modify the classes, and the embedded dependency is then re-signed by $A_{1}$ once the process completes, as in any vendoring process. The resulting uber-JAR is distributed by $A_{1}$, who is responsible for the packaged software. The Embedder ignores signature files and copies all other files without modification. We refrain from altering non-class files. If these other files are used, they must be referenced by other class files — which are themselves embedded. If they are not used, the Java Virtual Machine will not load them, and hence we don’t need to get provenance later. After processing each dependency JAR, it repackages it into a new JAR. It saves it in a directory, which we refer to as the augmented package repository, as shown in Figure 1.
@@
3.1.1 Embedder as a Maven plugin
@@
Within Classport, we create the custom embed goal. We configure it to fulfill the following needs of the plugin: 1. to have project’s class files available locally. This is fundamental because they have to be embedded. 2. To have dependencies’ class files. Dependencies must be resolved to also have both direct and transitive ones.
The first point is achieved by compiling the target project before executing the plugin. It guarantees that the class files for the project are available for processing. The second point is fulfilled by ensuring that the resolution of dependencies happens before the embedding process [^4] https://maven.apache.org/plugin-tools/apidocs/org/apache/maven/plugins/annotations/ResolutionScope.html. The embedding and then packaging process happens during a single Maven package phase.
3.1.2 Information embedded as Java Annotations
@@ 5 String version(); // V
The annotation contains the GAV coordinate, i.e., group, artifact, and version of the dependency to which the class belongs. The Retention annotation [19], @Retention(RetentionPolicy.RUNTIME), is used to make the ClassportInfo annotation visible at runtime. After the injection of Software Supply Chain information by the Embedder, the target application is packaged as a uber-JAR.
3.2 Introspector
@@
3.2.1 Introspector as Java agent using the Instrumentation API
@@ 11 Attributes:
1commons-io,commons-io,2.18.0
2commons-logging,commons-logging,1.3.4
3info.picocli,picocli,4.7.6
4org.apache.pdfbox,fontbox,3.0.4
5org.apache.pdfbox,pdfbox,3.0.4
6org.apache.pdfbox,pdfbox-examples,3.0.4
7org.apache.pdfbox,pdfbox-io,3.0.4
8org.apache.pdfbox,pdfbox-tools,3.0.4
9org.apache.pdfbox,preflight,3.0.4
10org.apache.pdfbox,xmpbox,3.0.4
11org.bouncycastle,bcpkix-jdk18on,1.80
12org.bouncycastle,bcprov-jdk18on,1.80
13org.bouncycastle,bcutil-jdk18on,1.80
3.4 Usage
The application developer is expected to run the packaging phase in Maven with our plugin configured. Our Maven plugin automatically embeds the dependency metadata into the compiled classes, producing an enriched JAR that can be published and distributed to package registries. Users of such JARs get the novel ability to introspect the embedded dependency tree at runtime using the Introspector component. Since Classport is automatically done within the packaging phase, keeping the embedded dependency information up-to-date is fully automated in the release process.
3.5 Implementation
Classport is implemented in Java. It is made of two modules: the Embedder uses the ASM [^6] https://asm.ow2.io/ library to manipulate the bytecode in the class files, while the Introspector is a Java agent that uses the Instrumentation API [20]. The implementation is open-source and available on GitHub. [^7] https://github.com/chains-project/classport
4 Experimental Methodology
@@ 4.1 Research Questions
Table 1: Study subjects considered in our experiments.
Project #Module Version Dependencies
PDFBox 13 3.0.4 19
Certificate Ripper 1 2.4.1 5
mcs 1 0.7.3 5
BioJava 14 7.2.4 50
checkstyle 1 10.23.0 37
GraphHopper 11 11.0 172
4.2 Dataset
We aim at evaluating Classport with a dataset of real-world, maintained, open-source Java projects, which use Maven and span a range of application domains, dependency sizes, and usage scenarios. Our dataset includes six projects, with projects varying in complexity, and includes both simple tools such as mcs and large frameworks such as Checkstyle. The most active repository is Checkstyle, which has around 9K stars on GitHub. Table 1 lists each considered Maven project along with its number of modules, version, and number of dependencies. The modules refer to the Maven modules (also called subprojects) of the project. For example, PDFBox has 13 modules and app and core are two of them. The considerred version is the tagged release of the project. The number of dependencies is calculated directly by Maven. We compile and embed all projects for RQ1 and analyze the resulting artifact. For RQ2, we run the Introspector while executing the tests of the projects in the dataset.
4.3 Methodology of RQ1
@@
Dependency Completeness This property ensures that Classport embeds the uber-JAR with dependency metadata of all dependencies. To assess this, we first run the Embedder on the dataset, generating the annotated uber-JARs. We only consider dependencies that are needed for the application’s execution, excluding test-only dependencies, as they do not end up in the uber-JAR. Then, we analyze the uber-JARs to read the annotations inside the binary artifacts. We need to only consider the uber-JAR because it contains all the dependencies (as opposed to a standard JAR which only contains the classes of the project and not the dependencies). We consider the embedding process complete in terms of dependency metadata if the dependency metadata, i.e., each Group ID, Artifact ID, and Version, matches the ground truth retrieved by running the Maven command mvn dependency:list on the project under test.
Class Completeness We assess completeness by verifying that the annotation is embedded in each class file of every dependency in the project. Although the dependency can have files other than class files, such as build manifests and various configuration files [21], we only consider class files for annotation as they are the only ones that can be executed. To verify completeness, we first run the Embedder on the study subjects. Then, we analyse the uber-JAR, looking inside each class file. If all Java class files are annotated with the Classport annotation, the embedding process is complete in terms of class metadata.
Build Time Overhead Build time overhead is measured by comparing the time taken to build the uber-JAR with and without the Embedder. In detail, we run package phase of the Maven build process with and without the Embedder and measure the time taken for each run using the Unix time command. We do the runs 10 times to accommodate for variance and we report the median of the overhead incurred.
@@
4.4 Methodology of RQ2
The second research question evaluates the behavior of Classport at runtime, in particular, its capability to introspect runtime dependencies. The objective is to show that 1) Classport can identify which dependencies are currently running and 2) the target application’s runtime behavior is not affected by Classport.
@@ 4.4 Methodology of RQ2
Runtime Dependency We report the dependencies detected during workload execution. As workload, we run the test suite of each study subject in Table 1 on the embedded project and record the runtime dependencies it triggers. This produces a list of runtime dependencies observed by Classport. The observed list should contains all embedded dependencies which are exercised at runtime.
Overhead To measure the runtime impact of Classport, we measure the time taken to execute the test suite with and without the Classport agent. We also do this measurement 10 times and report the median of the overhead incurred.